I need to find the slop of any line perpendicular to the line through points (0,5) and (-3,-4)
Answer by tutorcecilia(2144)

(0,5) and (-3,-4)
Lines that are perpendicular have slopes that are the negative reciprocals.
For example if:
Slope of line #1 = 1/2
Slope of Line #2=-2/1
Than the lines are perpendicular
.
First find the slope of the given pair of points:
Slope=m=(y1-y2)/(x1-x2)
Plug-in the values:
Slope=m=(5-(-4))/(0-(-3))=9/3=3
So,
Slope of Line #1=3=3/1
Slope of Line #2=-1/3 (the negative recirpocal of the slope of line #1)